# Personalized Generative Low-light Image Denoising and Enhancement

### v1

@@ Abstract
While smartphone cameras today can produce astonishingly good photos, their performance in low light is still not completely satisfactory because of the fundamental limits in photon shot noise and sensor read noise. Generative image restoration methods have demonstrated promising results compared to traditional methods, but they suffer from hallucinatory content generation when the signal-to-noise ratio (SNR) is low. Recognizing the availability of personalized photo galleries on users’ smartphones, we propose Personalized Generative Denoising (PGD) by building a diffusion model customized for different users. Our core innovation is an identity-consistent physical buffer that extracts the physical attributes of the person from the gallery. This ID-consistent physical buffer provides a strong prior that can be integrated with the diffusion model to restore the degraded images, without the need of fine-tuning. Over a wide range of low-light testing scenarios, we show that PGD achieves superior image denoising and enhancement performance compared to existing diffusion-based denoising approaches.
Figure 1: Using gallery photos from a user’s smartphone, we can restore low-light, noisy facial images. Our method produces finer details and better identity compared to current state-of-the-art face restoration approaches like DR2 [53] and GFP-GAN [50].
@@ 1 Introduction
where $\mathbf{x}$ is the unknown clean image, $\alpha$ is the sensor gain, $\eta$ is the dark current, and $\epsilon$ is the read noise. The inverse problem associated with low-light denoising is to recover $\mathbf{x}$ from $\mathbf{y}$. Over the past five decades, the main driving force is to exploit image structures, from Wiener filter [35], total variation [43, 3], wavelets [8] to non-local means [2], BM3D [9], and recently convolution neural networks [24, 61, 48, 5, 1, 17, 55], transformers [34, 37, 60], and diffusion models [25, 58, 66]. A common theme across existing methods is that they all focus on a generic prior $p(\mathbf{x})$, i.e., a prior learned from a large collection of example images. While they perform well in moderately difficult problems, the generality of these methods cannot be extended to heavily corrupted images. Particularly in the context of human face recovery, the restored results often lack real identity and exhibit artifacts. This is attributed to the ill-posed nature of the problem and the lack of appropriate constraints.
Why Gallery Photos? Smartphone cameras today store hundreds if not thousands of a user’s photos, captured at different times, in different places, and under different lighting conditions. While these images have many variations, they are all about the same person(s). Therefore, if the imaging goal is to take a photo of this user, the gallery on the phone would be the best source to build a prior $p(\mathbf{x})$. The situation is summarized in Fig. 2. In the context of diffusion-based image restoration, the original solution space can be large because many candidate solutions are consistent with the noisy observation. The gallery provides a strong constraint to the search problem. This allows us to search for better quality images with high precision of the person’s identity.
@@
  • What class of images to focus on: In this paper, we will focus on human faces because of their relevance to most smartphone users although our algorithm can be equally applied to other image classes. We assume that the gallery photos have been selected and are reasonably informative. Product-level engineering such as pre-processing and selection of the gallery photos is important but beyond the scope of this paper.
  • How to incorporate this class-specific $p(\mathbf{x})$ into the restoration model: The bigger (and harder) question is that given the gallery photos, how do we efficiently extract the prior information and improve the restoration? We assume that the gallery set is not too big and so we cannot train a restoration model from scratch. We want to avoid fine-tuning as much as possible during inference. Moreover, we want to make sure that the person’s identity is preserved.
Physical Buffers to the Rescue. Given the gallery photos, what kind of prior information would be useful for restoration? Advancements in computer vision have made it possible to extract detailed facial physical buffers—including albedo and normal maps—from a person’s gallery of images[42, 63]. These physical buffers capture intrinsic properties such as skin color and surface geometry, effectively encoding an individual’s unique identity and fine facial details. (See Fig. 3.) At the same time, it also eliminates the influence of environmental lighting, pose, and other identity-independent variables. We will leverage this rich prior information to improve restoration.
To convince readers that physical buffers could be useful to our denoising task, we show in Fig. 3 a toy experiment. We consider three types of degradation: low-resolution (LR), motion blur, and noise. In each case, we run an existing physical buffer extraction method DECA[14] to extract the albedo and surface normal of the input. We observe that despite the different types of degradation, the physical buffers (albedo and surface normal) are fairly accurate. The only failure case is when the noise is extremely strong. Therefore, if we can develop solutions to manage the heavy noise case, we will be able to use physical buffers as a signal prior to guiding the denoising process.
@@
    Personalized Low-Light Denoise and Enhancement Framework: We introduce a new framework that leverages diffusion models and person-specific priors from photo galleries to address identity loss in restored low-light facial images.
@@
    Easy Deployment: Our method operates easily without requiring fine-tuning for individual users, making it practical for real-world applications and scalable to multiple users.
@@ 2 Related Works
Figure 3: For images affected by mild low resolution, motion blur, or noise, DECA [14] can accurately extract physical buffers. However, DECA fails to obtain precise physical buffers for images degraded by severe low light and noise.
@@
Reference Prior for Face Restoration. In face restoration, reference images [46, 12, 31, 32] provide essential priors for recovering detailed and high-frequency facial features. These priors guide the model to generate realistic restorations. Reference-based methods can be divided into two types: single reference and gallery-based images. The single reference approach [46] uses one image to guide the restoration, focusing on enhancing facial details from a specific reference. In contrast, the gallery-based approach [12] leverages multiple reference images, offering more diverse information and stronger constraints, improving the accuracy and realism of the restoration.
@@ 3.1 Restoration by Diffusion
A diffusion model approximates the clean image distribution $p_{\theta}(x_{0})$ by learning a model $\theta$ that reverses the noise-adding process. In Denoising Diffusion Probabilistic Models (DDPMs) [20], Gaussian noise is progressively added to a clean image $x_{0}$:
@@
$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon,\quad\text{where}\quad\epsilon\sim\mathcal{N}(0,I)$ (2)
@@
The reverse process then removes the noise to reconstruct the clean image. After training, for any given time $t$ and the corresponding noisy image $x_{t}$, the model can iteratively denoise by sampling from $p(x_{0}|x_{t})$.
@@
In restoration tasks, like low-light denoising, the goal is to recover a high-quality image $x_{0}$ from a low-quality input $y_{0}$. Unlike pure generation, where the output can be any image that follows the distribution of clean images, restoration is constrained by the content of $y_{0}$. During training, the model uses $y_{0}$ as a condition, and once the model is trained, for any given time $t$ and the corresponding noisy image $x_{t}$, it can iteratively denoise by sampling from $p(x_{0}|x_{t},y_{0})$ using the trained model.
Figure 4: Path to the idea. Using facial physical buffers and gallery images can effectively preserve human face identity information.
@@
Figure 5: The overall architecture of the proposed method. Our core idea is to use ID-consistent physical buffers, extracted from gallery photos, to constrain the generative space in the diffusion model restoration process. For a high-quality gallery, we use LAP [64] to extract the albedo and normal information for each photo and apply adaptive aggregation to fuse the entire gallery. The extracted albedo represents base skin color and texture details, such as skin tone, freckles, and pigmentation, while normal captures facial geometry, including wrinkles, pores, and fine surface variations. In our framework, the output physical buffers isolate the intrinsic ID properties from lighting, shading, and pose, enabling the diffusion model to apply only ID-related information consistently.
3.2 Why Gallery + Physical Buffer?
@@
We formalize our core approach of extracting identity-consistent physical buffers from gallery photos to guide the diffusion model for restoration. This method is based on the observation that using facial physical buffers or gallery images can effectively preserve identity information during restoration tasks involving diffusion models. To evaluate the capability of physical buffers and gallery images in preserving identity information without relying on prior identity knowledge from degraded inputs, we train the following models using a blind scheme, where the degraded input ($y_{0}$) is not used as a condition during training. As illustrated in Fig. 4, the models are designed as follows:
  • Model 1: Trained exclusively on generic images as a purely generative model.
@@
  • Model 2: Trained on generic images while conditioned on the facial physical buffer extracted from the input image. This model incorporates face identity and facial features through the physical buffer.
  • Model 3: Fine-tuned on Model 1 with the user’s gallery photos. This model learns identity features from the target person’s gallery photos during fine-tuning.
@@
  • Model 4: Fine-tuned on Model 2 with the user’s gallery photos. This model integrates facial identity and features from the physical buffer of the input image, as well as the target person’s gallery photos.
We test these models using a simple low-light noisy input. Model 1 generates random faces, as expected for a purely generative model. In contrast, Model 2 produces images with facial features (e.g., nose, mouth, skin color) resembling those of the user. This is because the physical buffers extracted from the degraded user image (refer to Fig. 3) were used as conditioning. These results demonstrate that facial features and identity information can be incorporated by leveraging physical buffers. The outputs from Model 3 are all different user faces, compared to Model 1, showing that fine-tuning with gallery photos can effectively transfer strong identity information. The outputs from Model 4 are the best-expected results, demonstrating the effectiveness of using gallery photos and physical buffers to maintain the person’s identity.
However, despite the powerful combination of gallery images and physical buffer, there are limitations:
@@
    Fine-tuning: Extracting identity information from gallery photos via fine-tuning is time-consuming and inconvenient. Each time when we switch to a different user, the model must be fine-tuned again. Additionally, since the size of a gallery is usually much smaller than a pre-trained model, fine-tuning can easily overfit.
@@
    Extracting physical buffers from the degraded input image: If the input is heavily degraded, the physical buffers extracted from it can be wrong (refer to Fig. 3), and result in terrible results as shown in Fig. 4. Besides, the facial feature extracted from a single image is limited, not to mention from a degraded image.
In conclusion, we want to avoid fine-tuning the model on the gallery and we do not want to extract incorrect physical buffers due to noise.
3.3 ID-Consistent Physical Buffer
To address these limitations, we propose extracting physical buffers directly from existing clean gallery photos and using them as conditions to guide the diffusion model during training. This approach offers significant benefits: clean gallery photos allow us to form comprehensive and robust identity-consistent physical buffers containing detailed facial identity features of the target person. Specifically, We will extract the albedo and normal maps of the target person, which are two key attributes used to describe the physical characteristics of a face.
  • Albedo captures the intrinsic color of an object’s surface—that is, the subject’s color properties without any lighting effects.
  • Normal information is used to describe the fine bumps and details such as skin wrinkles and pores.
By conditioning on these strong identity physical buffers, we eliminate the need to fine-tune the model with gallery images and avoid the issue of incorrect buffer extraction. Consequently, the model is better equipped to preserve the target person’s identity and facial features, even under severe degradation.
3.4 Extracting Physical Buffers
Extracting physical buffers for a person from a set of gallery photos is a more challenging task than extracting these buffers from a single input, as in methods like DECA. This complexity arises because gallery photos of the same person are often taken under varying conditions, such as different scenes and poses. To ensure these buffers comprehensively represent the person’s identity and facial details, they must be aligned to maintain identity consistency. Fortunately, existing works have addressed this challenge. As shown in the physical buffer extractor module in Fig. 5, we adopt the aggregation net design from the LAP [64] to extract the ID-consistent physical buffers, which adaptively combines facial features from multiple images to learn consistent geometry and texture representations, and generates ID-consistent albedo and normal maps from a photo collection of the same individual.
@@ 3.5 Model Framework
Within the aggregation network, it contains a shared encoder $\delta$ across multiple images and a global decoder $\phi$ for predicting consistent face. To model albedo and normal, two separate aggregation networks denoted as ($\delta^{a},\phi^{a}$) and ($\delta^{n},\phi^{n}$), are employed. Given a photo gallery of N images $\{\textbf{I}_{k}^{i}\}_{k=1}^{N}$ of person i, each image is processed by $\delta^{a}$ and $\delta^{n}$ to extract its texture and geometry latent codes $x_{k}^{a}$ and $x_{k}^{n}$, respectively. To derive a global representation of the identity for person i based on $\{x_{k}^{a},x_{k}^{n}\}_{k=1}^{N}$, an adaptive aggregation method is adopted, This method learns a weight set $\{w_{k}^{a},w_{k}^{n}\}_{k=1}^{N}$ according to the quality of each image in$\{\textbf{I}_{k}^{i}\}_{k=1}^{N}$. The combined global ID-code $x_{c}^{a}$ and $x_{c}^{n}$ for texture and depth are calculated as:
$x_{c}^{a}=\sum_{k=1}^{N}w_{k}^{a}x_{k}^{a},\quad x_{c}^{n}=\sum_{k=1}^{N}w_{k}^{n}x_{k}^{n}.$ (3)
Then the global ID-code $x_{c}^{a}$ and $x_{c}^{n}$ are then passed to the decoders $\phi^{a}$ and $\phi^{n}$, respectively, to produce the ID-consistent albedo $a^{i}$ and normal $n^{i}$.
Additionally, with the low-light noisy image $y_{0}^{i}$ as another condition during training, other features, such as pose, hairstyle, and background inside the image beyond facial features can be effectively captured as well.
After the model is trained, during testing, the process begins with a Gaussian noise image $x_{T}$, which is iteratively transformed into the clean, enhanced image $x_{0}^{i}$ by sampling from $p(x_{0}^{i}|x_{T},y_{0}^{i},a_{i},n_{i})$ using the trained model. Our experiments demonstrate that identity information is effectively preserved with the guidance of these ID-consistent buffers, enabling our model to achieve superior results in human face restoration.

### v2

@@ Abstract
Modern cameras produce remarkably high-quality images, yet their performance in low-light conditions remains suboptimal due to fundamental limitations in photon shot noise and sensor read noise. While generative image restoration methods have shown promising results compared to traditional approaches, they often suffer from hallucinatory content generation when the signal-to-noise ratio (SNR) is low. Leveraging the availability of personalized photo galleries on users’ smartphones, we introduce Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that builds a customized diffusion model for individual users. Our key innovation lies in the development of an identity-consistent physical buffer that extracts the physical attributes of the person from the gallery. This ID-consistent physical buffer serves as a robust prior that can be seamlessly integrated into the diffusion model to restore degraded images without the need for fine-tuning. Over a wide range of low-light testing scenarios, we show that DiffPGD achieves superior image denoising and enhancement performance compared to existing diffusion-based denoising approaches.
Figure 1: Using gallery from a phone user, we can restore low-light noisy facial images. Our method produces finer details and better identity consistency compared to other state-of-the-art restoration approaches like MIRNet [61, 62], GFP-GAN [52], DiffLL [22] and CodeFormer [67] on both simulated (top row) and real-captured (bottom row) low-light noisy inputs.
@@ 1 Introduction
where $\mathbf{x}$ is the unknown clean image, $\alpha$ is the sensor gain, $\eta$ is the dark current, and $\epsilon$ is the read noise. The inverse problem associated with low-light denoising is to recover $\mathbf{x}$ from $\mathbf{y}$. Over the past five decades, the main driving force is to exploit image structures, from Wiener filter [36], total variation [44, 4], wavelets [9] to non-local means [3], BM3D [10], and recently convolution neural networks [25, 64, 50, 6, 2, 18, 57], transformers [35, 38, 63], and diffusion models [26, 60, 68]. A common theme across existing methods is to find and utilize a generic prior $p(\mathbf{x})$, i.e., a prior learned from a large collection of example images. While they perform well in moderately difficult problems, the generality of these methods cannot be extended to heavily corrupted images. Particularly in the context of human face recovery, the restored results often lack real identity and exhibit artifacts. This is attributed to the ill-posed nature of the problem and the lack of appropriate constraints.
Why Gallery Photos? Smartphone cameras today typically store hundreds if not thousands of a user’s photos, captured at different times, in different places, and under different lighting conditions. While these images have many variations, they are all about the same person(s). Therefore, if the imaging goal is to take a photo of this user, the gallery on the phone would be the best source to build a prior $p(\mathbf{x})$. This concept is illustrated in Fig. 2. In the context of diffusion-based image restoration, the initial solution space can be large since many candidate solutions are consistent with the noisy observation. The gallery provides a strong constraint to the search problem. This allows us to search for better quality images and preserve the person’s identity.
@@
  • What class of images we should focus on: In this paper, we focus on human faces because of their relevance to phone users. We assume that the gallery photos have been selected and are reasonably informative. Product-level engineering such as pre-processing and selection of the gallery photos is important but beyond the scope of this paper.
  • How to incorporate the class-specific $p(\mathbf{x})$ into the restoration model: The bigger (and harder) question is that given the gallery photos, how do we efficiently extract the prior information and improve the restoration? We do not want to train a restoration model from the scratch because the gallery can be small. We want to avoid fine-tuning as much as possible during inference. Moreover, the person’s identity needs to be preserved.
Physical Buffers to the Rescue. Given the gallery photos, what kind of prior information would be useful for restoration? Advancements in computer vision have made it possible to extract facial physical buffers—including albedo and normal maps—from a person’s gallery of images[43, 65, 47]. These physical buffers capture crucial identity-defining properties such as surface geometry and skin color, effectively encoding an individual’s unique identity. At the same time, it also eliminates the external influence of environmental lighting, pose, and other identity-independent variables. By incorporating this robust prior that disentangles the external influences, our approach can effectively extracts identity information from the gallery.
@@
    Personalized Low-Light Denoise and Enhancement Framework: We introduce a new framework that leverages diffusion models and person-specific priors from photo galleries to denoise low-light facial images while presering the identity.
@@
    Easy Deployment: Our method operates easily without requiring fine-tuning for individual users, making it practical and scalable to multiple users.
@@ 2 Related Works
Figure 3: Path to the idea. Using facial physical buffers and gallery images can help preserve human face identity information.
@@
Reference Prior for Face Restoration. In face restoration, reference images [48, 13, 32, 33] provide essential priors for recovering detailed and high-frequency facial features. These priors guide the model to generate realistic restorations. Reference-based methods can be divided into two types: single reference and gallery-based images. The single reference approach [48] uses one image to guide the restoration, focusing on enhancing facial details from a specific reference. In contrast, the gallery-based approach [13] leverages multiple reference images, offering more diverse information and stronger constraints, improving the accuracy and realism of the restoration. Despite the effectiveness [13] using gallery photos of users, it still requires multi-stage training and fine-tuning on each user during inference.
@@ 3 Method
Figure 4: For images affected by mild low-light noise, phycial buffers can be accurately extracted. However, when encountering the images degraded by severe low light and noise, precise physical buffers are difficult to obtained.
@@ 3.1 Restoration by Diffusion
A diffusion model approximates the clean image distribution $p_{\theta}(\mathbf{x}_{0})$ by learning a model $\theta$ that reverses the noise-adding process. In Denoising Diffusion Probabilistic Models (DDPMs) [21], Gaussian noise is progressively added to a clean image $\mathbf{x}_{0}$:
@@
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\text{where}\quad\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (2)
@@
The reverse process then removes the noise to reconstruct the clean image. After training, for any given time $t$ and the corresponding noisy image $\mathbf{x}_{t}$, the model can iteratively denoise by sampling from $p(\mathbf{x}_{0}|\mathbf{x}_{t})$.
@@
In restoration tasks, like low-light denoising, the goal is to recover a high-quality image $\mathbf{x}_{0}$ from a low-quality input $\mathbf{y}$. Unlike pure generation, where the output can be any image that follows the distribution of clean images, restoration is constrained by the content of $\mathbf{y}$. During training, the model uses $\mathbf{y}$ as a condition, and once the model is trained, for any given time $t$ and the corresponding noisy image $\mathbf{x}_{t}$, it can iteratively denoise by sampling from $p(\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y})$ using the trained model.
@@
Figure 5: The overall architecture of the proposed method. Our core idea is to use ID-consistent physical buffers, extracted from gallery photos, to constrain the generative space in the diffusion model restoration process. For a high-quality gallery, we use LAP [65] to extract the albedo and normal information for each photo and apply adaptive aggregation to fuse the entire gallery. The extracted albedo represents base skin color and facial appearance, while normal captures facial geometry. In our framework, the output physical buffers isolate the intrinsic ID properties from lighting, shading, and pose, enabling the diffusion model to apply only ID-related information consistently.
3.2 Why Combining Gallery and Physical Buffer?
@@
We formalize our core approach of extracting identity-consistent physical buffers from gallery photos to guide the diffusion model for restoration. This method is based on the observation that using facial physical buffers or gallery images can help preserve identity information during restoration tasks involving diffusion models. To evaluate the capability of physical buffers and gallery images, we conducted four simple experiments under a blind scheme (excluding $y_{0}$ during training), as illustrated in Fig. 3:
  • Exp 1: Model trained solely on generic images (purely generative).
@@
  • Exp 2: Model trained on generic images, conditioned on the facial physical buffer extracted from the input image.
  • Exp 3: Model trained on generic images, then fine-tuned with the user’s gallery photos.
@@
  • Exp 4: Model trained on generic images, fine-tuned with the user’s gallery photos, and conditioned on the facial physical buffer.
We tested these models on a mildly noisy input. Exp 1 produces random faces, as expected from a purely generative model. In contrast, Exp 2 uses physical buffers obtained via DECA [15], which can extract coarse face buffers from a single 2D photo (see Fig. 4), yielding faces with user-like features (e.g., nose shape, mouth and skin tone). Hence, physical buffers can convey key identity cues. Exp 3 outputs distinct faces of the same user, compared to Exp1, showing that fine-tuning with gallery photos effectively transfers identity information. Finally, Exp 4 yields the best results, showcasing that the collaborative use of gallery photos and facial physical prior can achieve the optimal outcomes.
While Fig. 3 shows us a design direction, we need to overcome:
@@
    Fine-tuning: Extracting identity via fine-tuning requires significant time and must be repeated for each user. Additionally, the small size of a gallery compared with a large pre-trained model can lead to overfitting.
@@
    Unreliable buffers from degraded inputs: As studied in the face reconstruction literature, low-quality inputs consistently present additional challenges [15]. Severe degradation undermines the extraction of physical buffers (Fig. 4), resulting in poor reconstruction. Moreover, features captured from a single degraded photo are limited, not to mention from a degraded one.
In conclusion, to avoid repeated fine-tuning on galleries and reduce the risk of erroneous buffers from noisy inputs, it is more desirable to jointly employing gallery photos and facial physical priors, as we will discuss below.
3.3 Extracting Physical Buffer from Photo Gallery
Based on the preceding analysis, we propose extracting physical buffers directly from existing clean gallery photos and using them as conditions to guide the diffusion model during training. This approach offers significant benefits: clean gallery photos allow us to form comprehensive and robust identity-consistent physical buffers containing more acurate facial identity features of the target person. Specifically, We will extract the albedo and normal maps of the target person, which are two key attributes used to describe the physical characteristics of a face.
  • Albedo captures the intrinsic color of an object’s surface—that is, the subject’s color properties without any lighting effects. It reveals a person’s facial appearance, e.g. eyebrows, lip coloration, and other distinct texture elements can reflect individual identity traits.
  • Normal encodes the surface orientation at each pixel, which provides hints about the underlying shape characteristics.
By conditioning on these identity physical buffers, we eliminate the need to fine-tune the model with gallery images and avoid the issue of incorrect buffer extraction. Consequently, the model is better equipped to preserve the target person’s identity and facial features, even under severe degradation.
However, extracting physical buffers for a person from a set of gallery photos is a more challenging task than extracting these buffers from a single input, as in methods like DECA. This complexity arises because gallery photos of the same person are often taken under varying conditions, such as different scenes and poses. To ensure these buffers comprehensively represent the person’s identity and facial properties, they must be aligned to maintain identity consistency. As shown in the physical buffer extractor module in Fig. 5(a), we adopt the aggregation net design from the LAP [65] to extract the ID-consistent physical buffers, which adaptively combines facial features from multiple images to learn consistent geometry and texture representations, and generates ID-consistent albedo and normal maps from a photo collection of the same individual.
@@ 3.4 Model Framework
The aggregation network contains a shared encoder $\delta$ across multiple images and a global decoder $\phi$ for predicting the consistent representation. To model albedo and normal, two separate aggregation networks, denoted as ($\delta^{a},\phi^{a}$) and ($\delta^{n},\phi^{n}$), are employed. Given a photo gallery of N images $\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$ of person i, each image is processed by $\delta^{a}$ and $\delta^{n}$ to extract its texture and geometry latent codes $\mathbf{\alpha}_{k}^{a}$ and $\mathbf{\alpha}_{k}^{n}$, respectively. To derive a global representation of the identity for person i based on $\{\mathbf{\alpha}_{k}^{a},\mathbf{\alpha}_{k}^{n}\}_{k=1}^{N}$, an adaptive aggregation method is adopted, This method learns a weight set $\{\mathbf{w}_{k}^{a},\mathbf{w}_{k}^{n}\}_{k=1}^{N}$ according to the quality of each image in$\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$. The combined global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ for texture and depth are calculated as:
$\mathbf{\alpha}_{c}^{a}=\sum_{k=1}^{N}\mathbf{w}_{k}^{a}\mathbf{\alpha}_{k}^{a},\text{ and}\quad\mathbf{\alpha}_{c}^{n}=\sum_{k=1}^{N}\mathbf{w}_{k}^{n}\mathbf{\alpha}_{k}^{n}.$ (3)
Once calculated, the global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ are passed to the decoders $\phi^{a}$ and $\phi^{n}$, respectively, to produce the ID-consistent albedo $\mathbf{a}^{i}$ and normal $\mathbf{n}^{i}$.
During inference, the process begins with a Gaussian noise image $\mathbf{x}_{T}^{i}$, which is iteratively transformed into the clean, enhanced image $\mathbf{x}_{0}^{i}$ by sampling from $p(\mathbf{x}_{0}^{i}|\mathbf{x}_{T}^{i},\mathbf{y}^{i},\mathbf{a}^{i},\mathbf{n}^{i})$ using the trained model. With the low-light noisy image $\mathbf{y}^{i}$ as another condition during training, other features, such as pose, hairstyle, and background inside the image beyond facial features can be effectively captured as well.
Our experiments demonstrate that identity information is effectively preserved with the guidance of these ID-consistent buffers, enabling our model to achieve superior results in human face restoration.